A moose is pulling a heavy sled through the snow (with friction). The sled has a mass of 65 kg. The moose applies a force of 780 N to the sled. The coefficient of kinetic friction between the sled and the snow is 0.26.
1
Expert's answer
2020-11-17T13:02:01-0500

Answer

Friction force="\mu mg\\=0.26\times65\times9.81\\=165.78N"

Applied force(F) =780N


Acceleration can be given as

"F-f=ma\\a=\frac{780-165.78}{65}\\a=9.45m/s^2"